# CFCPalsy: Facial Image Synthesis with Cross-Fusion Cycle Diffusion Model for Facial Paralysis Individuals

### v3

@@ 1 Introduction
In recent years, machine learning (ML) has shown promise in aiding the diagnosis of facial paralysis by analyzing facial features and expressions [63, 1, 15, 16]. However, the development of reliable ML models is hindered by the scarcity of comprehensive, annotated datasets that accurately represent the diversity and severity of facial paralysis. Obtaining high-quality facial images of facial palsy patients is a significant challenge due to several factors, including the need to protect patient privacy and the inherent sensitivity of medical data. Additionally, the complexity of the condition and the high labor costs involved make it challenging to collect a diverse and representative dataset. Existing datasets [65, 18, 25] are often limited in size, scope, and variability, restricting the generalization ability and performance of ML algorithms.
@@
Several studies have explored the use of generative adversarial networks (GANs) [17] to generate facial paralysis face images [67, 51, 50]. Although [67] successfully utilizes GANs to simulate facial expressions by reconstructing facial images from extracted boundary information, the generated images often lack the fine details necessary for accurate medical diagnosis, especially when there is a significant disparity between the facial features of the patient and the face to be combined with the palsy features. Building on this, another study [51] introduces facial shape normalization to address the unnatural appearance of images resulting from shape differences between the patient and the model. While this method improves image realism, it still struggles with capturing complex facial expressions and fine details, limiting its effectiveness in more nuanced diagnostic scenarios. In addition, to leverages data augmentation, GANs are also used to improve the performance of convolutional neural networks (CNNs) in grading facial palsy [50]. This approach synthesizes additional face images with varying degrees of facial palsy to combat overfitting and enhance the model’s generalization capability. However, the synthesized images in this study primarily focus on only five categories without adequately capturing the full range of facial expressions and subtle asymmetries inherent in real-world cases.
@@
In this paper, we aim to address the critical need for a comprehensive facial paralysis dataset by synthesizing a high-quality dataset. The generator is designed to capture a wide range of facial paralysis conditions and able to generate face images with varying degrees of facial paralysis based on a given facial identity image, providing a valuable resource for training ML models with the potential to improve diagnostic accuracy and clinical outcomes.
@@
To this end, we propose a Cross-Fusion Cycle Palsy Expression Generative Model (CFCPalsy). CFCPalsy is a unique triple condition image generation model, conditioned on identity information of a given ID image, expression information and landmark information of a given style image. It could generate a realistic image of the face in the ID image exhibiting the facial palsy characteristics of the face in the style image. CFCPalsy employs landmark features and the corresponding loss to ensure that the model could fully capture the nuanced facial deformations, particularly the detailed changes in facial features such as eyes, mouth, and overall asymmetry. When the model is fed multiple features simultaneously, it often encounters challenges such as feature redundancy and conflicting information, which can lead to degraded performance [36]. The redundancy among features can cause the model to overweight certain aspects, while conflicting signals may lead to inconsistent learning and prediction. To address these issues, CFCPalsy utilizes a novel cross-fusion feature integration method. This approach effectively harmonizes the information from multiple features, enabling the model to better leverage the complementary aspects of each feature, thereby significantly improving the overall performance and robustness of the model. Training diffusion models [20, 54] on limited datasets often leads to inefficiencies in learning and suboptimal model performance due to the scarcity of data. To address these challenges, we focused on improving the training efficiency and overall performance of diffusion models under constrained data conditions by modifying the training strategy. Our approach optimizes the learning process and proposes the cycle diffusion framework enabling the model to achieve superior results with a limited dataset.
@@ 2 Related Work
Face generation has become a prominent area of research within computer vision, driven by advancements in generative models such as Generative Adversarial Networks (GANs) [17], Variational Autoencoders (VAEs) [32, 48, 35, 55], and diffusion models [20, 54]. Among these, GANs have been extensively utilized for various tasks such as face generation and augumentation [9, 23, 37, 45, 59, 66]. Various methods have been developed to ensure that the generated face retains the original identity while allowing for other modifications such as image styles or background [60, 5, 68, 6, 53, 47, 11, 2]. To attain a better performance, researchers have recently turned to diffusion models [24, 69, 58], which offer a more stable and probabilistic framework for face image generation. DCFace [31] proposes a dual-condition diffusion model that generates synthetic face datasets, producing high-quality face images with varied identities and styles. Our work investigates how the severity of facial palsy and changes in facial expressions influence the synthesis of realistic images.
@@
Diffusion models [20, 54] have emerged as a robust alternative to GANs for a wide range of image generation tasks [12]. Unlike GANs, which rely on an adversarial framework, diffusion models gradually refine noisy inputs through a series of steps, leading to more stable and high-fidelity outputs. Several studies [21, 39, 52, 56, 27] have employed advanced sampling strategies that enhance both the efficiency and the output quality of diffusion models such as the latent space diffusion [49, 61, 14, 44, 41]. In addition, there are also studies [4, 29, 72, 28, 13, 33, 43, 57, 27, 30] focused on enhancing the models’ training efficiency by modifying the training strategies. We propose a novel cycle diffusion strategy that integrates the second loss setting designed to enhance model performance, particularly when training on limited datasets.
3 Method
In this section, we present our Cross-Fusion Cycle Palsy Expression Model (CFCPalsy) designed to synthesize high-quality facial paralysis images (see Fig. 2). First, we mention the necessary preliminary concepts about the diffusion models (sec 3.1). Next, we introduce the various feature extractors used in CFCPalsy which enhance the diffusion model’s ability to capture more facial details by accurately extracting multiple features (sec 3.2). In sec 3.3, we propose the cross feature fusion, which mitigates the model’s over-reliance on individual features and enhances its overall learning capability across all features by exchanging query vectors formed from different features. Finally, we devise the cycle training strategy that improves the training efficiency and performance of the diffusion model on limited datasets by modifying the training process and introducing a secondary loss function (sec 3.4).
3.1 Preliminary
@@
3.2 Feature Extraction Module
@@
3.3 Cross-Fusion Module
The attention mechanism [62] has become a fundamental component in many state-of-the-art models, particularly for tasks requiring the integration of information from different sources or across different parts of a data sequence. At its core, the attention mechanism allows a model to weigh the importance of different elements within a sequence, dynamically focusing on the most relevant parts to improve decision-making or feature extraction. The cross-attention mechanism extends this concept to operate between two distinct sequences. This facilitates the fusion of information from different sources, making it particularly useful for tasks that require integrating features from multiple modalities or input streams.
We aim for the model to fully utilize the information from identity, expression, and landmarks when predicting noise. Although directly concatenating the features from these three sources and injecting them into the model is a feasible approach, it does not yield satisfactory results. The generated images either lose the identity information from identity images $X_{ID}$ or fail to retain the facial expression details from original images $X_{0}$. Building on the concept of the cross-attention mechanism, we perform mutual interaction among multiple features, where the query sequence from one feature guides the extraction of information from the key-value sequence of another feature.
@@ 3.3 Cross-Fusion Module
3.4 Cycle Diffusion Strategy
@@
4 Experiments
4.1 Datasets
AFLFP:
The Annotated Facial Landmarks for Facial Palsy (AFLFP) [65] dataset is a comprehensive and laboratory-controlled collection designed to support the analysis of facial paralysis. It consists of facial images from 88 subjects, each participating in 16 different expression videos, such as brow raise, closed smile, gentle eye closure, and open smile. For each expression, the dataset captures keyframes representing four critical states: neutral, onset, a mid-state between onset and peak, and peak. In total, the AFLFP dataset contains 5,632 facial images, with 1,408 images corresponding to each key state and 64 images per subject.
MEEI:
The Massachusetts Eye and Ear Infirmary (MEEI) [18] dataset is a carefully curated collection focused on facial palsy, comprising both videos and high-resolution images. The dataset includes 60 videos from a diverse group of participants, consisting of 9 healthy subjects and 51 patients with varying degrees of facial paralysis. Each video captures 8 distinct facial movements, which are also provided as individual images, resulting in a total of 480 high-resolution images. The subjects in the MEEI dataset are categorized based on their eFACE [3] scores, which assess the severity of facial palsy.
4.2 Evaluation Metrics
@@
To evaluate the effectiveness of image generation from different perspectives, for the synthesized images, we compute Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [64], Learned Perceptual Image Patch Similarity (LPIPS) [70], Perceptual Loss (PL) [26], on unseen data. For this multi-condition synthesis task, we compute these metrics separately for identity images and facial paralysis expression images to simultaneously evaluate identity preservation and expression transfer.
Fréchet Inception Distance (FID) [19] is widely used to evaluate the quality of image generation. It measures the distance between the distribution of generated images and real images by comparing the statistics of feature representations extracted from a pre-trained Inception network. A lower FID score indicates that the generated images are more similar to the real images, both in terms of visual quality and diversity. However, some studies [8, 46] have shown that inconsistencies in the class distribution between the generated dataset and the real dataset can affect the calculation of the FID score.
To address this issue, we propose the aligned FID (aFID) (See Fig. 4) to evaluate the quality of the multi-condition diffusion model. Specifically, for each facial palsy patient in the test set, we first use images of other individuals as identity images and the patient’s condition image as the facial palsy expression image. This process synthesizes an image where the patient’s facial palsy expression is transferred onto another individual. We then use this generated dataset as the facial palsy expression images and the original facial palsy patient’s images as identity images to generate a second set of synthetic images. Finally, we calculate the FID score between this generated dataset and the real dataset. Ideally, the images in these two datasets should be identical.
@@
4.3 Implementation details
Table 1: Quantitative results on the test set.
@@ 4.3 Implementation details
GS = Grid Size, LM = Facial Landmark Features, CF = Cross-Fusion Module, CD = Cycle Diffuion Strategy.
@@
4.4 Ablation Study
To evaluate the contribution of each component in our model, we conducted a series of ablation studies. These experiments assess the impact of removing or altering key modules on the model’s performance. The visualization results of the ablation study are shown in Fig 5. The quantitative results are shown in Tab. 1.
@@ 4.4 Ablation Study
To investigate the role of facial landmark features in image generation, we remove the facial landmark condition from CFCPalsy. The model no longer learns landmark features, and in the fusion module, only identity features and expression features are used for information exchange through query key interactions and concatenation. During training, the model also stops using the landmark loss for weight updates. The results in Tab. 1 show a significant increase in aFID, and the pixel-level comparison metrics for facial palsy expression also exhibit a clear decline. As seen in Fig. 5, the model is largely unable to correctly learn facial palsy expressions. This highlights the importance of facial landmark features in synthesizing detailed facial palsy expressions.
@@
Tab. 1 compares the performance of the model with and without the cross-fusion module. It can be observed that directly concatenating multiple features and injecting them into the model leads to a sharp increase in aFID by 10.66, making it the component with the most significant impact on overall performance. Combined with other metrics, this shows that the cross-fusion module plays a crucial role in improving both identity preservation and facial palsy feature transfer. This feature fusion module effectively enhances the model’s ability to learn from multiple features comprehensively.
@@
The design of Cycle Diffusion is intended to improve the training efficiency of the model on limited datasets. We conducted comparative experiments to verify the effectiveness of this strategy. As shown in Tab. 1, the cycle training strategy provides a noticeable performance improvement for the model. We tracked the training process and plotted the changes in aFID, with the results shown in Fig. 6. Compared to the standard diffusion model training strategy, Cycle Diffusion achieves higher training efficiency.
4.5 Comparison with Previous Methods
We compared CFCPalsy with a similar facial synthesis task, DCFace [31], using 3x3 and 5x5 grid sizes on unseen images. The visual comparison results and numerical comparison results are shown in Fig. 5 and Tab. 1, respectively. Although DCFace 5x5 achieves better scores in several generated data and facial palsy expression data metrics, this does not necessarily indicate a higher degree of facial palsy feature transfer. Pixel-level and perception-based metrics not only focus on facial expressions but also consider factors like lighting and background, which are not primary considerations in the task of facial palsy synthesis. Compared to DCFace 5x5, aFID of images generated by CFCPalsy decreases by approximately 22%, and CFCPalsy also demonstrates a clear advantage in terms of identity feature preservation.
5 Conclusion

### v4

@@ Abstract
Index Terms:
Diffusion Models, Facial Image Synthesis, Deep Learning, Facial Paralysis Images, Expression Transfer
@@ I Introduction
In recent years, machine learning (ML) has shown promise in aiding the diagnosis of facial paralysis by analyzing facial features and expressions [3, 4, 5, 1]. However, the development of reliable ML models is limited by the lack of comprehensive, annotated datasets representing the diversity and severity of facial paralysis. Collecting high-quality images of facial palsy patients is challenging due to privacy concerns and the sensitive nature of medical data. Existing datasets [6, 7, 8] are often limited in size, scope, and variability, restricting the generalization ability and performance of ML algorithms.
@@
Several studies have explored the use of generative adversarial networks (GANs) [9] to generate facial paralysis face images [10, 11, 12]. Although [10] effectively employs GANs to simulate facial expressions by reconstructing images from boundary data, the generated images often lack the fine-grained details required for accurate medical diagnosis, especially when significant differences exist between the patient’s facial features and the target face incorporating palsy characteristics. Building on this, [11] incorporates facial shape normalization to mitigate unnatural appearances caused by shape differences between patients and models. Although this approach enhances image realism, it remains limited in capturing intricate facial expressions and fine details, reducing its utility in nuanced diagnostic cases. In addition to employing data augmentation, GANs are utilized to enhance CNN performance in grading facial palsy [12]. By generating face images with varying degrees of palsy, this method mitigates overfitting and improves model generalization. However, the synthesized images are limited to five categories, failing to capture the full spectrum of facial expressions and subtle asymmetries observed in real-world cases.
@@
This paper addresses the critical need for a comprehensive facial paralysis dataset by synthesizing high-quality images. Our generator is designed to capture diverse paralysis conditions, producing face images with varying severity levels based on a given facial identity. This dataset serves as a valuable resource for training ML models, with the potential to enhance diagnostic accuracy and improve clinical outcomes.
@@
To this end, we propose the Cross-Fusion Cycle Palsy Expression Generative Model (CFCPalsy), a novel triple-condition image generation framework. CFCPalsy generates realistic facial images by combining the identity of a given ID image with the expression and landmark features of a style image. By leveraging landmark features and a corresponding loss function, the model effectively captures nuanced facial deformations, including detailed changes in features such as the eyes, mouth, and overall asymmetry, ensuring high fidelity in depicting facial palsy characteristics. When fed multiple features simultaneously, models often face challenges like feature redundancy and conflicting information, leading to degraded performance [13]. Redundancy can cause the model to overweight certain features, while conflicts result in inconsistent learning and predictions. To overcome these issues, CFCPalsy employs a novel cross-fusion feature integration method. This approach harmonizes information from multiple features, leveraging their complementary strengths to enhance the model’s performance and robustness significantly. Training diffusion models [14, 15] on limited datasets often results in inefficiencies and suboptimal performance due to data scarcity. To address this, we enhanced training efficiency and performance under constrained data conditions by introducing a modified training strategy. Our approach includes the cycle diffusion framework, which optimizes the learning process and enables the model to deliver superior results even with limited data.
@@ II Related Work
Face generation has become a prominent research area in computer vision, driven by advancements in generative models such as Generative Adversarial Networks (GANs) [9], Variational Autoencoders (VAEs) [16, 17], and diffusion models [14, 15]. GANs have been widely applied to tasks like face generation and augmentation [18, 19, 20], with methods developed to retain original identity while enabling modifications such as image styles or backgrounds [21, 22, 23]. Recently, researchers have turned to diffusion models [24, 25, 26], which provide a more stable and probabilistic framework for face image generation. For instance, DCFace [27] introduces a dual-condition diffusion model to generate synthetic face datasets with varied identities and styles. Building on this, our work explores how the severity of facial palsy and changes in facial expressions impact the synthesis of realistic images.
@@
Diffusion models [14, 15] have emerged as a powerful alternative to GANs for image generation tasks [28]. Unlike GANs, which rely on adversarial training, diffusion models iteratively refine noisy inputs, producing stable and high-fidelity outputs. Recent advancements include enhanced sampling strategies [29, 30, 31] and latent space diffusion techniques [32, 33, 34], which improve efficiency and output quality. Additionally, studies [35, 36, 37] have optimized training efficiency through modified strategies. Building on this, we propose a novel cycle diffusion strategy incorporating a secondary loss function to improve performance, particularly on limited datasets.
III Method
In this section, we introduce the Cross-Fusion Cycle Palsy Expression Model (CFCPalsy) for synthesizing high-quality facial paralysis images (see Fig. 2). Section III-A outlines the foundational concepts of diffusion models. In Section III-B, we describe the feature extractors integrated into CFCPalsy, which enhance the model’s ability to capture fine facial details by accurately extracting diverse features. Section III-C details the proposed cross-feature fusion mechanism, which reduces over-reliance on individual features by exchanging query vectors derived from different features, thereby improving overall learning capability. Lastly, Section III-D introduces a cycle training strategy that boosts training efficiency and performance on limited datasets through process modifications and the inclusion of a secondary loss function.
III-A Preliminary
@@
III-B Feature Extraction Module
@@
III-C Cross-Fusion Module
The attention mechanism [42] has become essential in many state-of-the-art models, particularly for tasks requiring integration across data sequences or sources. By dynamically weighting elements within a sequence, it focuses on the most relevant parts to enhance feature extraction and decision-making. Extending this, the cross-attention mechanism operates between two sequences, enabling effective fusion of information from multiple modalities or input streams.
We aim to enable the model to fully leverage identity, expression, and landmark information when predicting noise. While directly concatenating features from these sources and injecting them into the model is a straightforward approach, it often leads to unsatisfactory results—generated images either lose identity details from $X_{ID}$ or fail to preserve the facial expression features from $X_{0}$. To address this, we adopt a cross-attention mechanism, allowing mutual interaction among features. Here, the query sequence of one feature guides the extraction of information from the key-value sequence of another, ensuring a more effective integration of features.
@@ III-C Cross-Fusion Module
III-D Cycle Diffusion Strategy
@@
IV Experiments
IV-A Datasets
AFLFP
The Annotated Facial Landmarks for Facial Palsy (AFLFP) [6] dataset is a comprehensive, laboratory-controlled collection designed for the analysis of facial paralysis. It includes facial images from 88 subjects, each participating in 16 expression videos, such as brow raise, closed smile, gentle eye closure, and open smile. For each expression, the dataset captures keyframes corresponding to four critical states: neutral, onset, mid-state (between onset and peak), and peak. In total, the AFLFP dataset contains 5,632 facial images, with 1,408 images for each key state and 64 images per subject.
MEEI
The Massachusetts Eye and Ear Infirmary (MEEI) [7] dataset is a curated collection focused on facial palsy, comprising videos and high-resolution images. It includes 60 videos from 9 healthy subjects and 51 patients with varying degrees of facial paralysis. Each video captures 8 distinct facial movements, which are also provided as individual images, resulting in a total of 480 high-resolution images. Subjects are categorized based on their eFACE [43] scores, which assess the severity of facial palsy.
@@ IV-A Datasets
IV-B Evaluation Metrics
@@
To evaluate the effectiveness of image generation from different perspectives, we compute Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [44], Learned Perceptual Image Patch Similarity (LPIPS) [45], and Perceptual Loss (PL) [46] on unseen data. For this multi-condition synthesis task, these metrics are computed separately for identity images and facial paralysis expression images, allowing simultaneous evaluation of identity preservation and expression transfer.
Fréchet Inception Distance (FID) [47] is a widely used metric for evaluating image generation quality. It quantifies the distance between the distributions of generated and real images by comparing feature statistics extracted from a pre-trained Inception network. A lower FID score indicates greater similarity between the generated and real images in terms of visual quality and diversity. However, some studies [48, 49] have shown that inconsistencies in class distributions between the generated and real datasets can affect the FID score calculation.
TABLE I: Quantitative results on the test set.
@@
GS = Grid Size, LM = Facial Landmark Features, CF = Cross-Fusion Module, CD = Cycle Diffusion Strategy.
To address this issue, we propose the aligned FID (aFID) (see Fig. 4) to evaluate the quality of the multi-condition diffusion model. Specifically, for each facial palsy patient in the test set, we use images of other individuals as identity images and the patient’s condition image as the facial palsy expression image. This process synthesizes an image where the patient’s facial palsy expression is transferred onto another individual. Next, we use this generated dataset as the facial palsy expression images and the original patient’s images as identity images to generate a second set of synthetic images. Finally, we calculate the FID score between this generated dataset and the real dataset. Ideally, the images in these two datasets should be identical.
IV-C Implementation details
@@
IV-D Ablation Study
To evaluate the contribution of each component in our model, we conducted a series of ablation studies. These experiments assess the impact of removing or altering key modules on the model’s performance. The visualization results of the ablation study are shown in Fig 5. The quantitative results are shown in Tab. I.
@@ IV-D Ablation Study
To investigate the role of facial landmark features in image generation, we remove the facial landmark condition from CFCPalsy. Without learning landmark features, the model relies solely on identity and expression features in the fusion module, using query-key interactions and concatenation for information exchange. Additionally, the landmark loss is excluded from weight updates during training. The results in Tab. I show a significant increase in aFID, while pixel-level comparison metrics for facial palsy expressions also decline notably. As illustrated in Fig. 5, the model fails to accurately learn facial palsy expressions, underscoring the critical role of facial landmark features in synthesizing detailed facial palsy expressions.
@@
Tab. I compares the performance of the model with and without the cross-fusion module. It can be observed that directly concatenating multiple features and injecting them into the model leads to a sharp increase in aFID by 10.66, making it the component with the most significant impact on overall performance. Combined with other metrics, this shows that the cross-fusion module plays a crucial role in improving both identity preservation and facial palsy feature transfer. This feature fusion module effectively enhances the model’s ability to learn from multiple features comprehensively.
@@
The Cycle Diffusion strategy is designed to enhance training efficiency on limited datasets. To validate its effectiveness, we conducted comparative experiments. As shown in Tab. I, the cycle training strategy significantly improves the model’s performance. Additionally, we tracked the training process and plotted the changes in aFID, as illustrated in Fig. 6. Compared to the standard diffusion model training strategy, Cycle Diffusion demonstrates superior training efficiency.
IV-E Comparison with Previous Methods
We compared CFCPalsy with DCFace [27] on a facial synthesis task using 3x3 and 5x5 grid sizes on unseen images. The visual and numerical comparison results are presented in Fig. 5 and Tab. I, respectively. While DCFace 5x5 achieves better scores on certain metrics for generated and facial palsy expression data, these scores do not necessarily reflect superior facial palsy feature transfer. Pixel-level and perception-based metrics also account for factors like lighting and background, which are not the primary focus of facial palsy synthesis. Compared to DCFace 5x5, CFCPalsy achieves a  22% reduction in aFID and demonstrates a notable advantage in identity feature preservation.
V Conclusion
@@
Acknowledgment
This work was supported by Shandong Provincial Natural Science Foundation under Grant ZR2024QF018.